# Contrasting Features of Parton Energy Loss in Heavy-ion Collisions at RHIC and the LHC

### v4

@@ Abstract
where $A$, $p_{0}$, and $n$ are free parameters in the fit.
@@
Figure 3 delineates $R_{AA}(p_{T})$ for charged hadrons in (a) 0–5% and (b) 30–40% Pb+Pb collisions at 2.76 TeV [31], and for charged pions in (c) 0–5% and (d) 30–40% Pb+Pb collisions at 5.02 TeV [26]. All the datasets exhibit upward trends for $p_{T}\gtrsim$ 5 GeV/$c$. When we apply the same fitting approach and fix $\alpha$ at 0.5, the resulting fit curves (dashed lines) adequately capture all the data points. When we take $\alpha$ as a free parameter (solid curve), the extracted $\alpha$ values are consistent with 0.5 within statistical uncertainties.
@@
We further investigate whether other final-state leading particles also exhibit these features. Figure 4 shows similar rising trends of $R_{AA}$ at higher $p_{T}$ for (a) $\pi^{0}$ and (b) $\eta$ mesons in 0–10% Pb+Pb at 2.76 TeV [32], for (c) prompt $J/\psi$ mesons in 0–100% Pb+Pb at 5.02 TeV [27], and for (d) prompt $D^{0}$ mesons in 0–10% Pb+Pb at 5.02 TeV [28]. The $\alpha$ values extracted for $\pi^{0}$, $\eta$, and $J/\psi$ mesons are consistent with 0.5 within the fitted statistical uncertainties. The fits to the prompt $D^{0}$ data seem to show some tension between the varied and fixed $\alpha$ values, but the difference is a less-than-$2\sigma$ effect. The curve with the fixed parameter ($\alpha=0.5$) does appear to agree with all the data points within the uncertainties. More precise measurements of $D^{0}$ $R_{AA}$ are required to better constrain the value of $\alpha$.
Figure 5 displays $R_{AA}(p_{T})$ for muons originating from (a) charm and (b) bottom hadrons in 0–10% Pb+Pb collisions at 5.02 TeV. In both cases, the fit curves with $\alpha$ = 0.5 align with all the data points within uncertainties. The $\alpha$ values extracted from the free-parameter fits exhibit a slight deviation from 0.5, with less than $1.5\sigma$ significance.
@@
We do, however, also investigate the relationship between energy loss and path length at LHC energies. Previous examinations of RHIC data have revealed that the deduced fractional energy loss, $\Delta p_{T}/p_{T}$, is a linear function of medium initial entropy density (quantified by $\frac{1}{S}\frac{dN}{dy}$) across different centrality intervals, despite significant variations in the path length for traversing partons [8]. This suggests that the initial collision density plays a more significant role in determining parton energy loss than the path length traversed by those partons. We apply the same analysis as adopted in Ref. [8] to the LHC data and discover similar outcomes, as shown in Fig. 6. Now that fractional energy loss varies with $p_{T}$ according to the LHC $R_{AA}$ data, we plot $\Delta p_{T}/p_{T}$ as a function of $\frac{1}{S}\frac{dN}{dy}$ at different $p_{T}$ values for charged hadrons in Pb+Pb collisions at 2.76 TeV and for charged pions in Pb+Pb collisions at 5.02 TeV. In each case for each $p_{T}$ regime, a clear linear trend emerges, and the linearity is especially strong for higher $p_{T}$ scales, where parton fragmentation dominates particle production [43].
@@
Our parton energy loss scenario predicts that the elliptic flow ($v_{2}$) of jets or high-$p_{T}$ leading particles in A+A collisions of intermediate centrality, if measured with minimal nonflow effects, will be close to zero and significantly smaller than values calculated assuming a static path-length dependence. Therefore, it is challenging to explain the large $v_{2}$ values observed for jets and leading particles at high $p_{T}$ [45]. Previous calculations of nuclear modification factor and anisotropic flow values for jets at RHIC and the LHC, however, do not prevent this possibility. In Ref. [46], a wide variety of energy loss models are shown to be able to describe available RHIC and LHC data within experimental uncertainty for both strong and weak path-length-dependent models. While this study indicates that the absence of path-length dependence is disfavored in the explored models inspired by quantum chromodynamics, it does not constrain the path-length-dependent parameter to be larger than zero [46]. The exact degree of path-length dependence of parton energy loss remains an outstanding physics issue, which should be addressed with more precise data. It is puzzling why the observed $v_{2}$ for jets or leading particles shows a weak $p_{T}$ dependence, while the magnitude of parton energy loss exhibits a strong $p_{T}$ dependence. There are some experimental details worthy of more investigation such as the subtraction scheme for the $v_{2}$-dependent background. The nonflow correlations in elliptic flow measurements for jets and high-$p_{T}$ particles present an important factor to be accounted for as well. These experimental concerns should be adequately addressed before the path length dependence of the parton energy loss may be definitively determined.

### v5

@@ Abstract
where $A$, $p_{0}$, and $n$ are free parameters in the fit. The fit results for each of these parameters and the $\chi^{2}/$ndf (number of degrees of freedom) for each fit are listed in Tables 1 and 2.
Table 1: Fit results for the $p_{T}$ spectra in $p$+$p$ collisions at $\sqrt{s_{NN}}=$ 2.76 TeV from Fig. 1 using Eq. (2).
Data Set $A$ $p_{0}$ $n$ $\chi^{2}/$ndf
charged particle 4097 0.883 7.13 0.52
$\pi^{0}\rightarrow\gamma\gamma$ 312.0 0.696 6.94 0.82
$\eta\rightarrow\gamma\gamma$ 0.888 1.305 7.47 0.30
Table 2: Fit results for the $p_{T}$ spectra in $p$+$p$ collisions at $\sqrt{s_{NN}}=$ 5.02 TeV from Fig. 1 using Eq. (2).
Data Set $A$ $p_{0}$ $n$ $\chi^{2}/$ndf
$\pi^{\pm}$ $2.94\times 10^{12}$ 0.895 6.97 0.82
prompt $J/\psi$ $6.33\times 10^{9}$ 0.987 7.08 0.46
prompt $D^{0}$ $4.46\times 10^{8}$ 1.759 6.47 0.07
$c\rightarrow\mu$ $5.41\times 10^{7}$ 0.957 6.25 0.13
$b\rightarrow\mu$ $4.40\times 10^{5}$ 1.699 6.40 0.73
@@
Table 3: Fit results for $R_{AA}(p_{T})$ in Pb+Pb collisions at $\sqrt{s_{NN}}=$ 2.76 TeV from Figs. 3 and 4 using Eq. (4).
Data Set $S_{0}$ $\alpha$ $\chi^{2}/$ndf
charged particle (0-5%) 1.56 0.42 0.18
charged particle (30-40%) 0.80 0.43 0.51
$\pi^{0}$ (0-10%) 1.44 0.50 0.11
$\eta$ (0-10%) 1.73 0.41 0.08
Table 4: Fit results for $R_{AA}(p_{T})$ in Pb+Pb collisions at $\sqrt{s_{NN}}=$ 5.02 TeV from Figs. 3, 4, and 5 using Eq. (4).
Data Set $S_{0}$ $\alpha$ $\chi^{2}/$ndf
$\pi^{\pm}$ (0-5%) 1.34 0.54 0.73
$\pi^{\pm}$ (30-40%) 0.75 0.54 0.37
prompt $J/\psi$ (0-100%) 0.76 0.54 0.30
prompt $D^{0}$ (0-10%) 2.36 0.27 0.24
$c\rightarrow\mu$ (0-10%) 0.48 0.83 0.20
$b\rightarrow\mu$ (0-10%) 0.96 0.60 0.04
Figure 3 delineates $R_{AA}(p_{T})$ for charged hadrons in (a) 0–5% and (b) 30–40% Pb+Pb collisions at 2.76 TeV [31], and for charged pions in (c) 0–5% and (d) 30–40% Pb+Pb collisions at 5.02 TeV [26]. All the datasets exhibit upward trends for $p_{T}\gtrsim$ 5 GeV/$c$. When we apply the same fitting approach and fix $\alpha$ at 0.5, the resulting fit curves (dashed lines) adequately capture all the data points. When we take $\alpha$ as a free parameter (solid curve), the extracted $\alpha$ values are consistent with 0.5 within statistical uncertainties. The fit results for all the $R_{AA}$ studies performed in this study and the corresponding $\chi^{2}$/ndf are listed in Tables 3 and 4.
@@
We further investigate whether other final-state leading particles also exhibit these features. Figure 4 shows similar rising trends of $R_{AA}$ at higher $p_{T}$ for (a) $\pi^{0}$ and (b) $\eta$ mesons in 0–10% Pb+Pb at 2.76 TeV [32], for (c) prompt $J/\psi$ mesons in 0–100% Pb+Pb at 5.02 TeV [27], and for (d) prompt $D^{0}$ mesons in 0–10% Pb+Pb at 5.02 TeV [28]. The $\alpha$ values extracted for $\pi^{0}$, $\eta$, and $J/\psi$ mesons are consistent with 0.5 within the fitted statistical uncertainties. The fits to the prompt $D^{0}$ data seem to show some tension between the varied and fixed $\alpha$ values, but the significance of the deviation is only $1.6\sigma$. The fixed-parameter fit with $\alpha=0.5$ agrees with nearly all the $D_{0}$ data points within uncertainties. Low precision beyond our $R_{AA}$ fit range in $p_{T}$ for RHIC data makes it difficult to determine whether the observed $\alpha=0.5$ versus 1 behavior would also occur for heavy flavor leading particles. However, we would expect this to remain the case. Reference [33] presents various transport model studies incorporating different energy loss mechanisms, which show good agreement with current RHIC and LHC data. Notably, the Duke model extends the standard Langevin equation by including medium-induced gluon radiation effects, successfully predicting $R_{AA}$ values consistent with observed $D$ meson suppression at the LHC. Future high-precision $D^{0}$ $R_{AA}$ measurements at high $p_{T}$ from both the LHC and RHIC will be crucial for further constraining the value of $\alpha$.
Figure 5 displays $R_{AA}(p_{T})$ for muons originating from (a) charm and (b) bottom hadrons in 0–10% Pb+Pb collisions at 5.02 TeV. In both cases, the fit curves with $\alpha$ = 0.5 align with all the data points within uncertainties. The $\alpha$ values extracted from the free-parameter fits exhibit a slight deviation from 0.5, with less than $1.5\sigma$ significance. Decay kinematics likely play a role in the smearing of the particle spectra for these cases, which may explain why the $\alpha$ values extracted for the muon data seem to trend above the 0.5 value we observe for the other data studied here.
@@
We do, however, also investigate the relationship between energy loss and path length at LHC energies. Previous examinations of RHIC data have revealed that the deduced fractional energy loss, $\Delta p_{T}/p_{T}$, is a linear function of medium initial entropy density (quantified by $\frac{1}{S}\frac{dN}{dy}$) across different centrality intervals, despite significant variations in the path length for traversing partons [8]. In contrast, when plotted as a function of path length (represented by various parameterizations using the number of participating nucleons $N_{\rm part}$), $\Delta p_{T}/p_{T}$ fails to exhibit a universal trend across different collision systems [8]. This suggests that the initial collision density plays a more significant role in determining parton energy loss than the path length traversed by those partons. We apply the same analysis as adopted in Ref. [8] to the LHC data and discover similar outcomes, as shown in Fig. 6. Now that fractional energy loss varies with $p_{T}$ according to the LHC $R_{AA}$ data, we plot $\Delta p_{T}/p_{T}$ as a function of $\frac{1}{S}\frac{dN}{dy}$ at different $p_{T}$ values for charged hadrons in Pb+Pb collisions at 2.76 TeV and for charged pions in Pb+Pb collisions at 5.02 TeV. In each case for each $p_{T}$ regime, a clear linear trend emerges, and the linearity is especially strong for higher $p_{T}$ scales, where parton fragmentation dominates particle production [44].
@@
Due to limited statistics across multiple centralities, it is challenging to perform a similar measurement for the heavy-flavor hadrons studied here. Recent theoretical models incorporating both collisional and radiative energy loss mechanisms predict a constant value of $\Delta p_{T}/p_{T}$ beyond $p_{T}\sim 5$ GeV/$c$ for charmed hadrons in various collision systems at LHC energies [46]. This would disagree with the rising $p_{T}$ trends observed in the $D^{0}$ and $J/\psi$ $R_{AA}$ data in this study which point to $p_{T}$-dependent fractional energy loss in LHC data, unlike RHIC data. Although these trends are less prominent in the two muons from heavy quark $R_{AA}$ fits, we expect this is due to the decay kinematics of those measurements which may obscure the expected rising trend. Leading hadron measurements provide a much cleaner measurement, thus we argue that the predicted flat fractional energy loss trends from these models undervalue the weight of radiative energy loss which would likely produce the observed rising trends. These discrepancies highlight the importance of future heavy-flavor $R_{AA}$ measurements at sPHENIX and the LHC to investigate these effects further.
Our parton energy loss scenario predicts that the elliptic flow ($v_{2}$) of jets or high-$p_{T}$ leading particles in A+A collisions of intermediate centrality, if measured with minimal nonflow effects, will be close to zero and significantly smaller than values calculated assuming a static path-length dependence. Therefore, it is challenging to explain the large $v_{2}$ values observed for jets and leading particles at high $p_{T}$ [47, 48]. Previous calculations of nuclear modification factor and anisotropic flow values for jets at RHIC and the LHC, however, do not prevent this possibility. In Ref. [49], a wide variety of energy loss models are shown to be able to describe available RHIC and LHC data within experimental uncertainty for both strong and weak path-length-dependent models. While this study indicates that the absence of path-length dependence is disfavored in the explored models inspired by quantum chromodynamics, it does not constrain the path-length-dependent parameter to be larger than zero [49]. The exact degree of path-length dependence of parton energy loss remains an outstanding physics issue, which should be addressed with more precise data. It is puzzling why the observed $v_{2}$ for jets or leading particles shows a weak $p_{T}$ dependence, while the magnitude of parton energy loss exhibits a strong $p_{T}$ dependence. There are some experimental details worthy of more investigation such as the subtraction scheme for the $v_{2}$-dependent background. The nonflow correlations in elliptic flow measurements for jets and high-$p_{T}$ particles present an important factor to be accounted for as well. These experimental concerns should be adequately addressed before the path length dependence of the parton energy loss may be definitively determined.